OpenStudy (anonymous):

why does (xy)^1/2 =sqrt of xy and not sqrt of x^2y^2

OpenStudy (amistre64):

Becasue 4^1/2 = 2, sqrt(4) = 2

OpenStudy (dumbcow):

n^1/2 = sqrt(n) in this case n=xy

OpenStudy (ricardovg):

(xy)^1/2 = x^(1/2)*y^(1/2)